OpenStudy (abbycross167):

36. A student scored an 88, 92, and 76 on three math tests. What does the student need to get on the fourth test to have an average of 85 for all four tests? Joseph wants to use algebra to solve this problem. Which equation should he use? (NOTE: The answers with the “4” or/and “3” under them are fractions) a. 4x = 256/85 b. 256 + x = 85 4 c. 85 + x = 25 4 d. 256 = 85x 3 4 My answer~d

OpenStudy (campbell_st):

well if x is the missing score you have 3 scores and the average of 4 scores needs to be 85 so the equation sets up as \[85 = \frac{88 + 92 + 76 + x}{4}\] so multiply both sides of the equation by 4 \[340 = x + 256\] now you can some for x

OpenStudy (abbycross167):

so does that mean it's b? or did I get it right?

OpenStudy (campbell_st):

well I'd say B is the solution \[\frac{256 + x}{4} = 85\]

OpenStudy (abbycross167):

ok, thank you soo much!!